USE THE EQUATION OF THE CURVE TO FIND THE ACCELERATION g
PLEASE HELP
Use the Cricket graph software and the Macintosh computer to graph the data Vave versus t. (vave is vertical axis t is horizontal axis.) curve of best fit. b) Find the equation ofthe curve of best fit. cl Use the equation of the curve to find the acceleration g Note: The quanuty that will give g is the coefficient or the t term. In addition, in order to convert this value from units of cm/0.017s/0.17s to units of cm/s this number must by multiplied by 60 X 6 360.
4. Calculate relative error in the value of g obtained in Calculations 2 and the accepted value of g at this latitude is g 981 cm/s2.


Time Velocity 0.017 0.034 0.051 0.068 0.085 0.102 0.119 0.136 0.153 0.17 0.187 0.204 0.221 0.238 0.255 0.272 0.289 0.306 0.323 0.34 35.29 35.29 41.18 64.71 82.35 94.12 100.00 135.29 147.06 152.94 170.59 194.12 205.88 217.65 235.29 523.53 282.35 305.88 323.53 329.41

Velocity v Time y 1044.8x 600 500 400 300 200 100 0 0.05 0.1 0.15 0.2 0.25 0.3 0.35 0.4 100
Answer #1

v= u+at

equation of motion of a particle in one dimensional

here u = 0 cm/s

a = g = acceleration due to gravity = slope of the linear fit

g = 1044.8 cm/s2

relative error = (1044.8 -981)/981 = 0.065 = 6.5%

please rate it up thanks :)